### v3

@@ 3.3 Oscillation mode frequency regimes
such that a frequency in the co-rotating frame, will be detected as by an observer. While this geometrical shift with may be relatively small for high-frequency modes of a slow rotator, the effect is large for low-frequency modes of a fast rotator. As an example, let us consider typical low-degree prograde () or retrograde () modes of Slowly Pulsating B (SPB) or Doradus (Dor) stars. These modes typically have frequencies of about 5 to Hz in a co-rotating frame of reference. These mode frequencies get shifted by in an observer’s inertial frame, as illustrated in Fig. 4. Following the measured rotation frequencies for these two classes of pulsators (cf. Fig. 6 in Aerts 2021), this geometrically based shift can reach values up to 30Hz and Hz for prograde and retrograde dipole () modes, respectively. Moreover, the frequency shifts increase dramatically as the mode degree increases, given that . In addition, the Coriolis force induces extra frequency shifts, lifting the degeneracy with respect to zonal () mode frequencies in the co-rotating frame. These shifts must additionally be taken into account, on top of the geometrical ones. It is then clear that one cannot make meaningful comparisons between observed frequencies as in Fig. 4 and their theoretical predictions in a co-rotating frame without applying the proper frequency shifts between the reference frames. This is an essential point of attention for asteroseismic modelling of moderate and fast rotators. Such modelling requires identification of the azimuthal orders in order to interpret the collective effect of all the numerous detected oscillations in observations of rotating stars whose modes and waves cannot be treated perturbatively with respect to the Coriolis force (cf. Fig. 3).
@@
In the following sections we discuss applications of asteroseismology with emphasis on moderate to fast rotators. However, we begin with slow rotators for ease of understanding and to focus on some recent highlights for such stars. We discuss applications for various types of stars, giving descriptions of case studies with the aim to encourage future applications.
@@ 4 Applications of asteroseismic modelling to single stars
In order to understand asteroseismology of fast rotators, it is convenient to first recall how it works for slow rotators. After a brief description of the basic principles of asteroseismic modelling, we remind the reader of some of the challenges involved in the probing of the Sun’s internal sound speed from its acoustic modes. Subsequently, we provide some recent topical updates of asteroseismology applied to slowly rotating low-mass stars before moving on to faster rotators. Throughout this section, Fig. 3 will be our guide for the applications. We start on the right with high-frequency acoustic oscillations and will gradually shift to the left along the frequency axis.
@@ 4.1 Principles of asteroseismic modelling of slow rotators
The linear free oscillation modes of a non-magnetic single star whose rotational flattening due to the centrifugal force can be ignored are obtained by solving the simplified version of Eq. (3), that is, by setting and ignoring all forces aside from the pressure force, gravity, and the Coriolis force. The wave equation then becomes
@@
where is the frequency in the co-rotating frame for simplicity of notation. The full expression of the linear operator is omitted here for brevity, as we will do for all theoretical expressions for the eigenmodes in the rest of this review. It can be found in Eq. 3.340 of Aerts et al. (2010). For modes in the super-inertial frequency regime, one can treat the Coriolis force in Eq. (13) perturbatively, while it has to be taken into account in full for moderate and fast rotators (cf. Fig. 3). In the applications spelled out below, we will gradually be upgrading towards higher complexity caused by rotation, focusing on pedagogy rather than completeness in quoting the literature.
@@
Simplifying maximally to zeroth order in (that is, no rotation) implies full separability of Eq. (13) in terms of spherical coordinates and time, as it reduces to Eq. (5). We thus find a family of spheroidal oscillation modes with frequencies . A degeneracy with respect to the azimuthal order occurs and the time-independent part for the Lagrangian displacement, , can be written in terms of spherical harmonics, . The expression for is derived in full details and spelled out explicitly in standard books treating non-radial oscillations of stars, see Unno et al. (1989) and Aerts et al. (2010, Chapter 3, Eq. 3.132).
@@
Forward asteroseismic modelling of slow rotators to estimate their basic stellar parameters such as mass, radius, core mass or envelope mass, and age is usually done by matching the observed frequencies, , of identified zonal () modes of degree and radial order , with those predicted from grids of 1D stellar evolution models. Such grid modelling is minimally a 4D optimisation problem, as any stellar evolution code will require the mass, initial chemical composition (any combination of the mass fractions of hydrogen , helium , or the metals ), and age as input, in order to compute the stellar structure for that moment in the star’s evolution. However, the problem to solve is in practice of much higher dimension, as numerous free parameters occur in the codes due to limitations in our knowledge of the input physics and/or simplifications of inherently 3D physical macroscopic processes into 1D prescriptions. While this 3D-to-1D simplification is fine for gravity, thermodynamics and the microphysics (nuclear reactions, equation-of-state, atomic diffusion, etc.), it gives bad approximations for macroscopic transport processes due to rotation, magnetism, and tides. Hence, for computations with codes assuming the stellar structure models to be described in 1D, one is forced to introduce free parameters summarising the effects of these 3D macroscopic phenomena. One thus rapidly ends up in a high-dimensional parameter space for the fitting of the oscillation frequencies, with ample opportunities nowadays to attack the regression problem with machine-learning tools (cf. Bellinger et al. 2016; Hendriks & Aerts 2019; Bellinger 2019, 2020; Angelou et al. 2020; Hon et al. 2020).
@@ 4.1.1 A few updates on solar modelling
The dimensionality of the regression problem involved in stellar modelling can be reduced appreciably if proper calibrations for the input physics or model-independent ranges for the parameters are available. We discuss this latter aspect for intermediate- and high mass stars in Sect. 5. For slowly rotating low-mass stars, one often freezes the input physics, notably the mixing length parameter used in time-independent convection theory, to parameter values calibrated from helioseismology. While doing so, it must be kept in mind that the models may include extra forces in the term in Eq. (1) that may be operational in the stars of the application, while being absent or less important for the Sun. We illustrate here that, even for the star we know best, helioseismology is still an active yet specialised sub-field of asteroseismology, for which the slow solar rotation matters in improving the dynamics in the Sun.
@@
In terms of physical ingredients, progress in solar models compared to the famous Model S (see Christensen-Dalsgaard et al. 1996, for details) is large (Christensen-Dalsgaard 2021). Yet, the physical description for two regions inside the Sun, notably the base of the convective envelope and the surface layers, require improvements. The properties of the internal rotation of the Sun as summarised by Thompson et al. (2003, Figure 7) and updated by Christensen-Dalsgaard & Thompson (2007) reveal a decrease in the surface rotation frequency from about 450 nHz at the solar pole (zero latitude) to about 330 nHz for latitude . The level of latitudinal differentiality decreases from the solar surface to the base of the convection zone situated at a fractional radius of . The transition region between the differentially rotating convective envelope and the quasi-rigidly rotating (at about 440 nHz) radiative interior is called the tachocline.
@@
Figure 5 illustrates the deviations between the squared sound speed profile of the true Sun delivered by its thousands of identified acoustic modes and those based on some of the best models. The agreement between the 1D solar models and helioseismic data is better than 1.5%. Yet, the remaining discrepancies between theory and observations are highly significant due to the extreme precision of the observations. Aside from the ongoing controversy between the ‘old’ and ‘new’ Sun in terms of its surface abundance determinations (Buldgen et al. 2023a, b), the deviations between the Sun and current solar models in the tachocline are likely linked to the interplay of processes at microscopic and macroscopic scales. This concerns limited knowledge in opacities (Bailey et al. 2015), in mixing processes at the base of the convection zone (Baraffe et al. 2022), and in angular momentum transport (Eggenberger et al. 2022a), among other aspects.
@@
All 1D models of the Sun and of low-mass stars also suffer from the so-called “surface effect” (see the lack of points between and 1 in Fig. 5). This represents an offset between measured frequencies of identified modes and those predicted by the best 1D solar models of order a few Hz, which is much larger than observational uncertainties. The offsets are mode dependent and increase for increasing radial order of the acoustic modes (higher mode frequencies). Improving envelope models of the Sun and of low-mass stars requires a full 3D description incorporating the interplay between acoustic oscillations and 3D time-dependent convection (Mosumgaard et al. 2020), keeping in mind the non-adiabaticity of the gas (Houdek et al. 2019).
@@
Novel 3D simulations of the whole convective solar envelope, including shear-driven magnetic buoyancy due to rotation in the tachocline, provide new insights into the local mixing it induces (Matilsky et al. 2022; Duguid et al. 2023). This illustrates that rotation matters, even for slow rotators such as the Sun. Also, thanks to long-term monitoring from space by the Heliospheric and Magnetic Imager onboard NASA’s Solar Dynamics Observatory (SDO, Pesnell et al. 2012), a novelty was introduced into helioseismology in 2018. Aside from the well-known spheroidal high-frequency acoustic modes occurring in the high-frequency regime to the right in Fig. 3 and used in Fig. 5, toroidal inertial modes restored by the Coriolis force with frequencies below twice the solar rotational frequency were discovered in the SDO data by Löptien et al. (2018). Those concern equatorial Rossby modes as well as high-latitude inertial modes (Gizon et al. 2021). These recently discovered retrograde modes open up inertial-mode helioseismology, an exciting new way to probe the dynamics of the rotating tachocline and solar envelope (Bekki et al. 2022). We come back to Rossby modes in Sect. 4.4, where it will be pointed out that such modes were already discovered in numerous rapidly rotating stars of intermediate mass years prior to those found in the Sun.
@@ 4.1.2 Updates on forward modelling of slow rotators
For the distant stars, we cannot rely on high-degree (say, ) modes from space photometry, as their effects cancel out when integrating the variability across the visible hemisphere in the line of sight. Forward modelling in such cases relies on observables in addition to the mode signals (Gent et al. 2022), notably spectroscopic and astrometric input to delineate the grids of 1D equilibrium models used for the regression in the modelling of the stellar interior. Nowadays, spectroscopic effective temperatures and gravities, as well as luminosities from Gaia parallaxes (Gaia Collaboration et al. 2016) have become available for large ensembles of low-mass dwarfs and red giants in the Milky Way. These inferred stellar properties are used in tandem with their seismic observables from CoRoT, Kepler and TESS for galactic studies (cf. Fig. 1, Anders et al. 2014; Chiappini et al. 2015; Pinsonneault et al. 2018; Zinn et al. 2019, 2020; Claytor et al. 2020; Mackereth et al. 2021; Grunblatt et al. 2021; Zinn et al. 2022). We do point out the large difference in relative precision between such “classical” quantities and the seismic observables connected with identified modes (Aerts et al. 2019, factors 10 to 1000, see Table 1 in ).
@@
The matching between observed and theoretically predicted frequencies of identified modes is mostly done in a Bayesian framework, where the error estimation is often tackled numerically from a Markov Chain Monte Carlo approach (Appourchaux et al. 2009; Gruberbauer et al. 2012, 2013; Chaplin et al. 2014; Silva Aguirre et al. 2017). Deducing stellar parameters with high precision for large ensembles is one major asset of asteroseismology (e.g., Stello et al. 2011; Hekker et al. 2011; Huber et al. 2011, 2012; Hekker et al. 2013; Davies et al. 2016; Brogaard et al. 2018; Farnir et al. 2021; Brogaard et al. 2022; Li et al. 2022b, 2023b, etc.) as input for other studies in astrophysics (cf. Fig. 1), notably galactic archaeology (Miglio et al. 2009, 2013; Stello et al. 2013; Ness et al. 2016; Stello et al. 2017; Anders et al. 2017b, a; Silva Aguirre et al. 2018; Hekker & Johnson 2019; Ness et al. 2019; Miglio et al. 2021; Ness et al. 2022; Li et al. 2022b; Hon et al. 2022; Anders et al. 2023, etc.), old open clusters (Brogaard et al. 2012; Miglio et al. 2012, 2016; Brogaard et al. 2021), and exoplanetary research where age-dating and star–planet interactions are crucial aspects (Huber et al. 2013; Lebreton & Goupil 2014; Silva Aguirre et al. 2015; Huber et al. 2019; Chontos et al. 2021; Huber et al. 2022). However, the ages of the stars deduced from any method, including asteroseismology, remain dependent on the input physics used to compute the stellar evolution models and their isochrones, no matter how precisely the mass and initial chemical composition of the stars have been derived. For this reason, stellar ages are at best precise, but not necessarily accurate.
@@
The unknown level of internal mixing due to element transport causes a major systematic uncertainty for the age-dating of stars (Salaris & Cassisi 2017), notably for those of intermediate or high mass born with a convective core. As an example, age-dating of red giants from non-rotating isochrones while ignoring the fast rotation of their progenitors on the main sequence may lead to systematic age uncertainties in the red giant phases of up to 20% (Fritzewski et al. 2024). However, this effect is often ignored when providing age estimates and their uncertainties for red giants. More generally, the cumulative effect of convective core overshooting during the main-sequence phase, which may partially be caused by rotational mixing and possibly be inhibited somewhat by a magnetic field, is among the most important unknown factors in age-dating, for all levels of internal rotation rates and masses (e.g. Deheuvels et al. 2015; Bellinger 2019; Mombarg et al. 2019; Pedersen et al. 2021; Johnston 2021; Noll et al. 2021; Noll & Deheuvels 2023).
@@
Developing a calibrated theory of transport processes (Mathis et al. 2013) for all stellar masses and across all evolutionary phases is thus a major global aim of asteroseismology. This is currently the only feasible method to offer a proper value for the numerous free parameters occurring in 1D implementations of these processes, because their parameters change on an evolutionary timescale. Major progress has been achieved to improve internal angular momentum transport, guided by the overarching measured properties of internal rotation across stellar evolution (Aerts et al. 2019). Simulations based on new theoretical ingredients, involving magnetic fields and/or internal gravity waves are successful in explaining the internal rotation rates deduced from high-precision space asteroseismology (Rogers et al. 2013; Fuller et al. 2014; Rogers 2015; Eggenberger et al. 2017, 2019; Fuller et al. 2019; Ratnasingam et al. 2020; Takahashi & Langer 2021; Eggenberger et al. 2022b; Moyano et al. 2023). Nevertheless, more work is needed when it comes to the understanding of asteroseismically determined levels of internal mixing for various ensembles of pulsators (Rogers & McElwaine 2017; Deal et al. 2018, 2020; Mombarg et al. 2020; Pedersen et al. 2021; Mombarg et al. 2022; Varghese et al. 2023). Asteroseismically measured values of envelope mixing assuming a diffusive coefficient for the element tranport in the stellar envelope range from 1 cm s to 10 cm s (Aerts 2021, Table 1). Both the level and functional form of internal shear mixing is the dominant unknown ingredient affecting the ages and convective core masses of stars across stellar evolution (Van Grootel et al. 2010a, b; Giammichele et al. 2018; Charpinet et al. 2019b; Tkachenko et al. 2020; Johnston 2021; Pedersen et al. 2021; Pedersen 2022b).
@@
Future age-dating of stars with 10% accuracy instead of precision from asteroseismically calibrated transport processes for exoplanet host stars is a challenging aim in the core science program of the ESA PLATO space mission (Rauer et al. 2014, and Rauer et al. 2024, submitted, for the latest update on the mission development). PLATO also offers an extensive Complementary Science program open to the worldwide community offering the study of internal rotation, magnetism, and mixing across the entire HRD (cf. Fig. 1).
@@
Making progress in our understanding of internal mixing due to element transport requires good knowledge of . This brings us back to the quest to deduce the internal rotation profile of stars for different phases of their evolution. Taking the Coriolis force in Eq. (13) into account lifts the degeneracy of the mode frequencies with respect to and gives rise to resolved rotationally split multiplets if the duration of the time series data is sufficiently long compared to the average rotation period of the star. Ignoring terms in in the equation of motion and adopting a perturbative 1st-order approach in for shellular rotation, leads to observed frequencies in an intertial frame of reference according to Ledoux splitting (Ledoux 1951) given by
@@
where are the so-called rotational kernels, which can be computed from the identified modes and the equilibrium structure of a 1D stellar model (Aerts 2021, Eq. 45). This expression includes the geometrical shift introduced in Eq. (18), as well as the Ledoux constant caused by the Coriolise force (Ledoux 1951). Hence, following Eq. (14), the rotation frequency throughout the stellar interior gives rise to frequency multiplets with components in the line-of-sight, provided that all the modes in the multiplet are excited to observable amplitude. Note that the latter circumstance is a good assumption for solar-like oscillation modes excited stochastically by envelope convection but that this is not necessarily the case for heat-driven oscillations, as shown in this era of high-precision space asteroseismology (Aerts 2021). Indeed, modes of particular may not be excited intrinsically by the heat mechanism, but may be pumped up in mode energy from non-linear mode interactions involving the rotation frequency. We give some examples of this below, as it has been measured in fast rotators.
@@
Equation (14) reveals that treating the Coriolis force perturbatively up to 1st-order creates symmetrical Ledoux splittings for fixed and , given that ranges from . Moreover, in the limit of high-order p and g modes, one can show that and , respectively (see Aerts et al. 2010, for a summary of derivations of these approximations). Hence, for such modes, the measured rotational splittings provide a direct estimate of the local rotational profile averaged by the mode energy, since is determined by the square of the mode displacement (cf. Eq. (45) in Aerts 2021).

### v4

@@ 3.3 Oscillation mode frequency regimes
such that a frequency in the co-rotating frame, will be detected as by an observer. While this geometrical shift with may be relatively small for high-frequency modes of a slow rotator, the effect is large for low-frequency modes of a fast rotator. As an example, let us consider typical low-degree prograde () or retrograde () modes of slowly pulsating B (SPB) or Doradus (Dor) stars. These modes typically have frequencies of about 5 to Hz in a co-rotating frame of reference. These mode frequencies get shifted by in an observer’s inertial frame, as illustrated in Fig. 4. Following the measured rotation frequencies for these two classes of pulsators (cf. Fig. 6 in Aerts 2021), this geometrically based shift can reach values up to 30Hz and Hz for prograde and retrograde dipole () modes, respectively. Moreover, the frequency shifts increase dramatically as the mode degree increases, given that . In addition, the Coriolis force induces extra frequency shifts, lifting the degeneracy with respect to zonal () mode frequencies in the co-rotating frame. These shifts must also be taken into account, in addition to the geometrical shifts. It is then clear that one cannot make meaningful comparisons between observed frequencies as in Fig. 4 and their theoretical predictions in a co-rotating frame without applying the proper frequency shifts between the reference frames. This is an essential point of attention for asteroseismic modelling of moderate and fast rotators. Such modelling requires the identification of the azimuthal orders in order to interpret the collective effect of all the numerous detected oscillations in observations of rotating stars whose modes and waves cannot be treated perturbatively with respect to the Coriolis force (see Fig. 3).
@@
In the following sections we discuss some applications of asteroseismology, with an emphasis on moderate to fast rotators. However, we begin with slow rotators for ease of understanding and to focus on some recent highlights for such stars. We discuss applications for various types of stars, giving descriptions of case studies with the aim to encourage future applications.
@@ 4 Applications of asteroseismic modelling to single stars
In order to understand the asteroseismology of fast rotators, it is convenient to first recall how it works for slow rotators. After a brief description of the basic principles of asteroseismic modelling, we recall some of the challenges involved in the probing of the Sun’s internal sound speed from its acoustic modes. Subsequently, we provide some recent topical updates of asteroseismology applied to slowly rotating low-mass stars before moving on to faster rotators. Throughout this section, Fig. 3 will be our guide for the applications. We start on the right with high-frequency acoustic oscillations and will gradually shift to the left along the frequency axis.
@@ 4.1 Principles of asteroseismic modelling of slow rotators
The linear free oscillation modes of a non-magnetic single star whose oblateness caused by the centrifugal force can be ignored are obtained by solving the simplified version of Eq. (3), that is, by setting and ignoring all forces aside from gravity and the pressure and Coriolis forces. The wave equation then becomes
@@
where is the frequency in the co-rotating frame for simplicity of notation. The full expression of the linear operator is omitted here for brevity, as we do for all theoretical expressions for the eigenmodes in the rest of this review. It can be found in Eq. 3.340 of Aerts et al. (2010). For modes in the super-inertial frequency regime, one can treat the Coriolis force in Eq. (13) perturbatively, while it has to be taken into account in full for moderate and fast rotators (see Fig. 3). In the applications spelled out below, we gradually upgrade towards higher complexity caused by rotation, focusing on pedagogy rather than completeness in quoting the literature.
@@
Simplifying maximally to zeroth order in (i.e. no rotation) implies the full separability of Eq. (13) in terms of spherical coordinates and time as it reduces to Eq. (5). We thus find a family of spheroidal oscillation modes with frequencies . A degeneracy with respect to the azimuthal order occurs, and the time-independent part for the Lagrangian displacement, , can be written in terms of spherical harmonics, . The expression for is derived in full detail and spelled out explicitly in standard books treating non-radial oscillations of stars (see Unno et al. (1989) and Aerts et al. 2010, Chapter 3, Eq. 3.132).
@@
Forward asteroseismic modelling of slow rotators to estimate their basic stellar parameters such as mass, radius, core mass or envelope mass, and age is usually done by matching the observed frequencies, , of identified zonal () modes of degree and radial order , with those predicted from grids of 1D stellar evolution models. Such grid modelling is minimally a 4D optimisation problem as any stellar evolution code requires the mass, initial chemical composition (any combination of the mass fractions of hydrogen , helium , or the metals ), and age as input in order to compute the stellar structure for that moment in the star’s evolution. However, the problem to solve is in practice of much higher dimension, as numerous free parameters occur in the codes due to limitations in our knowledge of the input physics and/or simplifications of inherently 3D physical macroscopic processes into 1D prescriptions. While this 3D-to-1D simplification is fine for gravity, thermodynamics, and the microphysics (e.g. nuclear reactions, equation-of-state, atomic diffusion), it gives bad approximations for macroscopic transport processes due to rotation, magnetism, and tides. Hence, for computations with codes assuming the stellar structure models to be described in 1D, one is forced to introduce free parameters summarising the effects of these 3D macroscopic phenomena. One thus rapidly ends up in a high-dimensional parameter space for the fitting of the oscillation frequencies; there are currently ample opportunities to attack the regression problem with machine-learning tools (cf. Bellinger et al. 2016; Hendriks & Aerts 2019; Bellinger 2019, 2020; Angelou et al. 2020; Hon et al. 2020).
@@ 4.1.1 A few updates on solar modelling
The dimensionality of the regression problem involved in stellar modelling can be reduced appreciably if proper calibrations for the input physics or model-independent ranges for the parameters are available. We discuss this latter aspect for intermediate- and high-mass stars in Sect. 5. For slow-rotating low-mass stars, one often freezes the input physics, notably the mixing length parameter used in time-independent convection theory, to parameter values calibrated from helioseismology. While doing so, it must be kept in mind that the models may include extra forces in the term in Eq. (1) that may be operational in the stars of the application, while being absent or less important for the Sun. We illustrate here that, even for the star we know best, helioseismology is still an active yet specialised sub-field of asteroseismology, for which the slow solar rotation matters in improving the dynamics in the Sun.
@@
In terms of physical ingredients, the progress in solar models compared to the famous Model S (see Christensen-Dalsgaard et al. 1996, for details) has been great (Christensen-Dalsgaard 2021). However, the physical description for two regions inside the Sun, notably the base of the convective envelope and the surface layers, require improvements. The properties of the internal rotation of the Sun as summarised by Thompson et al. (2003, Fig. 7) and updated by Christensen-Dalsgaard & Thompson (2007) reveal a decrease in the surface rotation frequency from about 450 nHz at the solar pole (zero latitude) to about 330 nHz at latitude . The level of latitudinal differentiality decreases from the solar surface to the base of the convection zone situated at a fractional radius of . The transition region between the differentially rotating convective envelope and the quasi-rigidly rotating (at about 440 nHz) radiative interior is called the tachocline.
@@
Figure 5 illustrates the deviations between the squared sound speed profile of the true Sun delivered by its thousands of identified acoustic modes and those based on some of the best models. The agreement between the 1D solar models and helioseismic data is better than 1.5%. However, the remaining discrepancies between theory and observations are highly significant due to the extreme precision of the observations. Aside from the ongoing controversy between the ‘old’ and ‘new’ Sun in terms of its surface abundance determinations (Buldgen et al. 2023, 2024), the deviations between the Sun and the current solar models in the tachocline are likely linked to the interplay of processes on microscopic and macroscopic scales. This is a question of limited knowledge of opacities (Bailey et al. 2015), of mixing processes at the base of the convection zone (Baraffe et al. 2022), and of angular momentum transport (Eggenberger et al. 2022a), among other aspects.
@@
All 1D models of the Sun and of low-mass stars also suffer from the what is known as the ‘surface effect’ (see the lack of points between and 1 in Fig. 5). This represents an offset between measured frequencies of identified modes and those predicted by the best 1D solar models of the order of a few Hz, which is much larger than observational uncertainties. The offsets are mode dependent and increase for increasing radial order of the acoustic modes (higher mode frequencies). Improving envelope models of the Sun and of low-mass stars requires a full 3D description incorporating the interplay between acoustic oscillations and 3D time-dependent convection (Mosumgaard et al. 2020), keeping in mind the non-adiabaticity of the gas (Houdek et al. 2019).
@@
Novel 3D simulations of the whole convective solar envelope, including shear-driven magnetic buoyancy due to rotation in the tachocline, provide new insights into the local mixing it induces (Matilsky et al. 2022; Duguid et al. 2023). This illustrates that rotation matters, even for slow rotators such as the Sun. In addition, thanks to long-term monitoring from space by the Heliospheric and Magnetic Imager on board NASA’s Solar Dynamics Observatory (SDO, Pesnell et al. 2012), a novelty was introduced into helioseismology in 2018. Aside from the well-known spheroidal high-frequency acoustic modes occurring in the high-frequency regime to the right in Fig. 3 and used in Fig. 5, toroidal inertial modes restored by the Coriolis force with frequencies below twice the solar rotational frequency were discovered in the SDO data by Löptien et al. (2018). They are related to equatorial Rossby modes as well as high-latitude inertial modes (Gizon et al. 2021). These recently discovered retrograde modes open up inertial-mode helioseismology, an exciting new way to probe the dynamics of the rotating tachocline and solar envelope (Bekki et al. 2022). We come back to Rossby modes in Sect. 4.4, where we point out that such modes were already discovered in numerous rapidly rotating stars of intermediate mass years prior to those found in the Sun.
@@ 4.1.2 Updates on forward modelling of slow rotators
For the distant stars, we cannot rely on high-degree (e.g. ) modes from space photometry, as their effects cancel out when integrating the variability across the visible hemisphere in the line of sight. Forward modelling in such cases relies on observables in addition to the mode signals (Gent et al. 2022), notably spectroscopic and astrometric input to delineate the grids of 1D equilibrium models used for the regression in the modelling of the stellar interior. Spectroscopic effective temperatures and gravities, as well as luminosities from Gaia parallaxes (Gaia Collaboration et al. 2016) have become available for large ensembles of low-mass dwarfs and red giants in the Milky Way. These inferred stellar properties are used in tandem with their seismic observables from CoRoT, Kepler and TESS for galactic studies (see Fig. 1; Anders et al. 2014; Chiappini et al. 2015; Pinsonneault et al. 2018; Zinn et al. 2019, 2020; Claytor et al. 2020; Mackereth et al. 2021; Grunblatt et al. 2021; Zinn et al. 2022). We note the large difference in relative precision between such classical quantities and the seismic observables connected with identified modes (Aerts et al. 2019, factors 10 to 1000, see Table 1 in ).
@@
The matching between observed and theoretically predicted frequencies of identified modes is mostly done in a Bayesian framework, where the error estimation is often tackled numerically from a Markov chain Monte Carlo approach (Appourchaux et al. 2009; Gruberbauer et al. 2012, 2013; Chaplin et al. 2014; Silva Aguirre et al. 2017). Deducing stellar parameters with high precision for large ensembles is one major asset of asteroseismology (e.g. Stello et al. 2011; Hekker et al. 2011; Huber et al. 2011, 2012; Hekker et al. 2013; Davies et al. 2016; Brogaard et al. 2018; Farnir et al. 2021; Brogaard et al. 2022; Li et al. 2022b, 2023b) as input for other studies in astrophysics (see Fig. 1), notably galactic archaeology (Miglio et al. 2009, 2013; Stello et al. 2013; Ness et al. 2016; Stello et al. 2017; Anders et al. 2017b, a; Silva Aguirre et al. 2018; Hekker & Johnson 2019; Ness et al. 2019; Miglio et al. 2021; Ness et al. 2022; Li et al. 2022b; Hon et al. 2022; Anders et al. 2023, etc.), old open clusters (Brogaard et al. 2012; Miglio et al. 2012, 2016; Brogaard et al. 2021), and exoplanetary research where age-dating and star–planet interactions are crucial aspects (Huber et al. 2013; Lebreton & Goupil 2014; Silva Aguirre et al. 2015; Huber et al. 2019; Chontos et al. 2021; Huber et al. 2022). However, the ages of the stars deduced from any method, including asteroseismology, remain dependent on the input physics used to compute the stellar evolution models and their isochrones, no matter how precisely the mass and initial chemical composition of the stars have been derived. For this reason, stellar ages are at best precise, but not necessarily accurate.
@@
The unknown level of internal mixing due to element transport causes a major systematic uncertainty for the age-dating of stars (Salaris & Cassisi 2017), notably for those of intermediate or high mass born with a convective core. As an example, the age-dating of red giants from non-rotating isochrones while ignoring the fast rotation of their progenitors on the main sequence may lead to systematic age uncertainties in the red giant phases of up to 20% (Fritzewski et al. 2024a). However, this effect is often ignored when providing age estimates and their uncertainties for red giants. More generally, the cumulative effect of convective core overshooting during the main-sequence phase, which may partially be caused by rotational mixing and possibly be inhibited somewhat by a magnetic field, is among the most important unknown factors in age-dating, for all levels of internal rotation rates and masses (e.g. Deheuvels et al. 2015; Bellinger 2019; Mombarg et al. 2019; Pedersen et al. 2021; Johnston 2021; Noll et al. 2021; Noll & Deheuvels 2023).
@@
Developing a calibrated theory of transport processes (Mathis et al. 2013) for all stellar masses and across all evolutionary phases is thus a major aim of asteroseismology. This is currently the only feasible method to offer a proper value for the numerous free parameters occurring in 1D implementations of these processes because their parameters change on an evolutionary timescale. Major progress has been achieved to improve internal angular momentum transport, guided by the overarching measured properties of internal rotation across stellar evolution (Aerts et al. 2019). Simulations based on new theoretical ingredients, involving magnetic fields and/or internal gravity waves are successful in explaining the internal rotation rates deduced from high-precision space asteroseismology (Rogers et al. 2013; Fuller et al. 2014; Rogers 2015; Eggenberger et al. 2017, 2019; Fuller et al. 2019; Ratnasingam et al. 2020; Takahashi & Langer 2021; Eggenberger et al. 2022b; Moyano et al. 2023). Nevertheless, more work is needed when it comes to the understanding of asteroseismically determined levels of internal mixing for various ensembles of pulsators (Rogers & McElwaine 2017; Deal et al. 2018, 2020; Mombarg et al. 2020; Pedersen et al. 2021; Mombarg et al. 2022; Varghese et al. 2023). Asteroseismically measured values of envelope mixing assuming a diffusive coefficient for the element transport in the stellar envelope range from 1 cm2 s-1 to 106 cm s-1 (Aerts 2021, Table 1). Both the level and functional form of internal shear mixing is the dominant unknown ingredient affecting the ages and convective core masses of stars across stellar evolution (Van Grootel et al. 2010a, b; Giammichele et al. 2018; Charpinet et al. 2019b; Tkachenko et al. 2020; Johnston 2021; Pedersen et al. 2021; Pedersen 2022b).
@@
Future age-dating of stars with 10% accuracy instead of precision from asteroseismically calibrated transport processes for exoplanet host stars is a challenging aim in the core science programme of the ESA PLATO space mission (Rauer et al. 2024). PLATO also offers an extensive Complementary Science programme open to the worldwide community offering the study of internal rotation, magnetism, and mixing across the entire HRD (see Fig. 1).
@@
Making progress in our understanding of internal mixing due to element transport requires a good knowledge of . This brings us back to the quest to deduce the internal rotation profile of stars for different phases of their evolution. Taking the Coriolis force in Eq. (13) into account lifts the degeneracy of the mode frequencies with respect to and gives rise to resolved rotationally split multiplets if the duration of the time series data is sufficiently long compared to the average rotation period of the star. Ignoring terms in in the equation of motion and adopting a perturbative first-order approach in for shellular rotation leads to observed frequencies in an intertial frame of reference according to Ledoux splitting (Ledoux 1951) given by
@@
where are the rotational kernels, which can be computed from the identified modes and the equilibrium structure of a 1D stellar model (Aerts 2021, Eq. 45). This expression includes the geometrical shift introduced in Eq. (18), as well as the Ledoux constant caused by the Coriolis force (Ledoux 1951). Hence, following Eq. (14), the rotation frequency throughout the stellar interior gives rise to frequency multiplets with components in the line of sight, provided that all the modes in the multiplet are excited to observable amplitude. We note that the latter circumstance is a good assumption for solar-like oscillation modes excited stochastically by envelope convection, but that this is not necessarily the case for heat-driven oscillations, as shown in this era of high-precision space asteroseismology (Aerts 2021). Modes of particular may not be excited intrinsically by the heat mechanism, but may be pumped up in mode energy from non-linear mode interactions involving the rotation frequency. We give some examples of this below, as measured in fast rotators.
@@
Equation (14) reveals that treating the Coriolis force perturbatively up to first-order creates symmetrical Ledoux splittings for fixed and , given that ranges from . Moreover, in the limit of high-order p and g modes, one can show that and , respectively (see Aerts et al. 2010, for a summary of derivations of these approximations). Hence, for such modes, the measured rotational splittings provide a direct estimate of the local rotational profile averaged by the mode energy, since is determined by the square of the mode displacement (see Eq. (45) in Aerts 2021).